Wouls some one please tell me how to do this? I am lost here. Thank you for your help.
Write 116 2/3% as a fraction or a mixed number

Answer by checkley71(8403)

116 2/3%=1.16 2/3=3*1.16+2)/3=5.48/3=548/300=274/150 ANSWER.